Consider the following numerical example of the simple keynesian model with no government spending ,taxes or foreign sector (all figures in R millions):
C=100 + 0,9Y
I=50
Answer the following question.
1- use a graph to illustrate the equilibrium level of output.
1
Expert's answer
2019-08-13T08:56:38-0400

The equilibrium level of output is:

Y = C+ I = 100 + 0.9Y + 50,

0.1Y= 150,

Y =1500.


The equilibrium level of output is shown as an intersection point of AD and AS(Y) curves.